(11square root of -12)(-12square root of -1): (11square root of -12)(-12square root of -1)
Answer by orca(336)

11sqrt(-12)(-12sqrt(-1))
=11sqrt((-1*12))(-12sqrt(-1)) ...................(-1 = i^2)
=11sqrt((3*2^2*i^2))(-12sqrt(i^2))
=11*2i*sqrt(3)(-12i)
=11*2*(-12)i^2sqrt(3)
=11*2*(-12)(-1)sqrt(3) ...................(i^2 = -1)
=264sqrt(3)